### v1

# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

Abstract
By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost. However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real world settings.
Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve model’s parameter efficiency. We validate the effectiveness of our method by pruning two state-of-the-art MoE models, Mixtral-8x7B and Mixtral-8x22B. Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. To facilitate future research, we will release our code and the pruned MoE models.
[…]
siunitx
Diversifying the Expert Knowledge for Task-Agnostic Pruning
in Sparse Mixture-of-Experts
Zeliang Zhang1111This research was conducted during the author’s internship at Microsoft Research.  Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
1 Introduction
Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws [16]. In addition to increasing the depth of neural network models, widening neural networks by using the sparsely-activated mixture-of-experts (MoE) architecture is also proved effective. MoE widens the feed-forward network (FFN) layer (one expert) by having multiple parallel FFNs (experts). During forward propagation, only a subset of these experts is activated. Thus, compared to dense models, MoE models achieve better end-task performance and generalize better to new tasks without increasing computation costs. Notable examples of MoE models include Switch Transformer [9], Mixtral-MoE [14], and Uni-MoE [22].
Despite significant progress in developing wider and deeper MoE LLMs, the increased memory consumption due to larger model sizes (i.e., increased number of experts) poses a substantial challenge to the deployment of these models in real-world settings. For example, storing and loading Mixtral-8x7B, which has 8 experts in each of its 32 layers, requires approximately $88$ GB. The MoE layers constitute the majority of the parameters. Adding or removing even one expert in each layer can significantly impact overall memory cost and model performance. For example,  [27] shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark [12]. In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
There are many recent studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. [3] prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. [5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. [21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. [11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. [27] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
Figure 1: Removing several experts from the original MoE layer would not cause model collapse but improve efficiency. Prior works often utilize expert access information to filter out unimportant experts. In our work, we first group different experts with similar knowledge in the feature space, then merge them along with the routers to prune the MoE layer. This post-processing approach allows us to diversify the features of each MoE layer, thereby preserving the knowledge of the original large models as much as possible while reducing computation and storage consumption.
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 1, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for two state of the art MoE models, Mixtral-8x7B and Mixtral-8x22B. Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
2 Related Work
Sparse MoEs. The Mixture-of-Experts (MoE) structure is firstly applied in classical machine learning models by Jacobs et al. [13] and Jordan and Jacobs [15], then widely used in various deep learning models [38, 30, 39]. Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models [33, 19, 42]. It adapts the original large feed-forward network (FFN) in each transformer block into multiple smaller FFNs, forming an expert layer with a router that computes the weighted output of each MoE layer. Sparse MoEs were first proposed by Fedus et al. [9]. In this approach, only a few experts are activated in each layer, accelerating training and inference while significantly increasing the number of parameters for greater model capacity. Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer [9], Mixtral-8B [14], and Uni-MoE [22]. Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models [40, 20].
Model Pruning. Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency [25, 36, 24]. Pruning techniques can be categorized into unstructured pruning [23, 35, 29], which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning [18, 8, 34], which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs [4]. Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs [28], RNNs [41], and LSTMs [7] to modern large models such as Llama [37] and Stable-Diffusion [2].
While the large amount of parameters in sparse MoEs benefits the model’s capacity to achieve good performance at the pre-training stage, the increasing memory consumption causes great challenges to fine-tuning different downstream tasks. In this paper, we work on pruning the sparse MoEs to reduce redundant experts in the task-agnostic setting, which enhances the computational and memory efficiency throughout the fine-tuning process, and scalability of deploying these models.
3 Methodology
Notations. We denote $F(\cdot;\Theta,W,K)$ as an MoE layer, where $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ indicates the parameter set for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, $W\in\mathbb{R}^{N\times d}$ indicates the routing weights for each expert and $K$ indicates the number of experts selected for each token.
For a new token $x\in\mathbb{R}^{d}$, we first compute the routing logits to score the match between $x$ and the $n$-th expert:
$p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}},$ (1)
where $W_{t}\in\mathbb{R}^{d}$ is the routing policy of $t$-th expert. Then, we take the experts $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$ with top-$K$ logits $\{l_{n}\}_{n=i_{1}}^{i_{K}}$ as the selected target experts. Last, we compute the weighted average of the outputs from selected experts as the final output of the MoE layer,
$y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$ (2)
3.1 Task Definition
Several prior works [27, 11, 21] have identified that removing some experts in different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
Specifically, for a MoE LLM $\mathcal{M}(\cdot;\mathcal{F})$, which consists of multiple MoE layers =$\mathcal{F}=\{F^{l}(\cdot;\Theta^{l},W^{l},K)\}_{l=1}^{L}$, we want to prune $r$ experts from the total $N$. This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete. We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ that satisfy the following optimization objective:
$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y),$ (3)
where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the pruned LLM by deleting $r$ experts from each MoE layer $F$ in $\mathcal{F}$, and $\mathcal{D}$ is the dataset.
This task is particularly challenging as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions. In this work, we propose to estimate expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
Figure 2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under two kernel-based CKA criteria (Linear and RBF). A darker color indicates a greater similarity between experts.
3.2 Expert Similarity in MoE layer
Motivation: Intuitively, the more independent the experts are within the same MoE layer, the more diverse the representations learned by each expert. This diversity makes it harder to replace or drop any of the experts. Conversely, the redundancy of shared knowledge learned by different experts provides opportunities for model pruning.
We employ the Centered Kernel Alignment (CKA) [17] as the criteria to evaluate the similarity between experts in each MoE layer. Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
$\rho_{ij}=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (4)
where $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 [32]. We respectively use the linear kernel and RBF kernel in similarity computation of eq. 4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
As illustrated in Fig. 2, the darker the cell, the more similar the two related experts. The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers. There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts. For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
Remark.
We notice that a concurrent work [26] also studies expert similarity, but focuses on the cosine similarity between expert weights. Our work examines the differences between experts using a more generalized metric based on the correlation between inputs and outputs, and obtains a different observation: even without considering the last layer, expert similarity does not gradually decrease with increasing depth of layers. For example, we observe higher pairwise expert similarity at the $11$-th layer compared to the $6$-th layer, which is cross-verified by both distance metrics. Additionally, the experts in the $26$-th layer also demonstrate a higher similarity with others compared with shallow layers. We provide the full evaluation results on visualizing the expert similarity of MoE layers in Mixtral-8x7B and Mixtral-8x22B in appendix A.
3.3 Discover and Merge Similar Experts on Feature Space
Figure 3: Overview of our method. We first leverage different metrics to obtain the expert representation, then compute the pairwise distance to get the expert similarity matrix. Based on the expert similarity matrix, we can group the experts with shared knowledge in the same cluster, which can be merged on the weight space for MoE pruning.
Having identified the existence of moderate expert similarity in each MoE layer, we propose pruning the experts within each MoE layer by merging experts with shared knowledge. We present an overview of our method in fig. 3. Suppose we want to reduce the number of experts from $N$ to $r$, the details of our approach are outlined as follows.
Discovering Similar Experts from the Expert Graph. We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ to group similar experts into the same clusters, where we denote $\mathcal{V}$ as the set of nodes and $\mathcal{E}$ as the connections between nodes.
First, for each expert $f_{i}$, we compute its feature representation $\mathcal{M}(f_{i})$, where $\mathcal{M}(\cdot)$ is a specialized function to represent $f_{i}$.
Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
Then, we initialize the $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
Last, we leverage the graph partition algorithm to split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$. The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ shares the most similar knowledge with each other, and shows much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$. To achieve this goal, we adopt a two-step strategy. At the first step, we partition the $\mathcal{V}$ of $\mathcal{G}$ into two subsets $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, where $\mathcal{V}_{1}$ has $r-1$ nodes and $\mathcal{V}_{2}$ has $G-r+1$ nodes. Then, we split $\mathcal{V}_{1}$ into $r-1$ subsets, $\mathcal{V}_{1,1},...,\mathcal{V}_{1,r-1}$. The splitting process can be formulated as follows,
$\min\sum_{i,j\in\mathcal{V}}x_{i}x_{j}D_{ij}-\sum_{i,j\in\mathcal{V}}x_{i}(1-x_{j})D_{ij}$ (5)
$-\sum_{i,j\in\mathcal{V}}(1-x_{i})(1-x_{j})D_{ij},$
[…]
where $x_{i}=1$ if node $i$ in group $\mathcal{V}_{1}$, and $x_{i}=0$ otherwise. The first term is used to minimize the distance between nodes in $\mathcal{V}_{1}$, the second term is usd to maximize the distance of nodes between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, and the third term is ued to maximize the distance of nodes in $\mathcal{V}_{2}$.
When pruning $G$ experts to $r$ experts, there are $r-1$ experts keeping unaffected in the entire process. Our graph partition strategy can preserve the most knowledge while diversifying the expert features in each MoE layers.
Merging Similar Experts with Routers. To preserve the different knowledge learned by different models clustered in the same group, we merge these experts as well as their routers on the weight space as follows,
$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},$ (6)
$\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
where $\mathcal{V}_{n}$ is the set of similar experts in the $n$-th cluster $\mathcal{G}_{n}$, and we have $\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}=1$. We update the MoE layer $F$ by respectively replacing all the experts grouped in $\mathcal{G}_{n}$ by a single FFN expert layer $f(\cdot;\hat{\theta}_{n})$ and corresponding routing weights $\hat{W}_{n}$, $n=1,2,...,N-r$.
3.4 Practical Implementation
While we have developed a general strategy to prune sparse MoEs, several practical questions arise during implementation:
Q1: How to decide $\mathcal{M}(\cdot)$ and perform task-agnostic pruning on sparse MoEs?
Intuitively, given the same input, similar experts should generate similar outputs. Thus, we propose to use the mean squared error on the outputs of two experts to evaluate the similarity between them.
We follow previous work [27] to select few samples from pre-trained dataset C4 [32] as the calibration dataset to compute the expert feature representation at the third step for similar experts discovery. The C4 dataset comprises multiple domains of contents, which can be a good surrogate to simulate the performance of the task-agnostic pruning.
However, with the limited size of the calibration dataset, the estimation of expert similarity based on the aforementioned expert representation could be inaccurate due to biased sampling. To improve the generalization ability at the discovery stage, we randomly mix up the token embedding for data augmentation and then use the augmented data for computing the expert representation and similarity. In practice, we iteratively update the expert similarity matrix by continuously shuffling and mixing up the data for full augmentation.
Q2: Is it possible to prune sparse MoEs without using any data?
Experts in the MoE layer of current models typically have the same structure; for instance, in Mixtral-8x7B, each expert is a two-layer FFN with SiLU activation. By training on different routed tokens, the experts usually learn different distributions during the training process. This results in a disparity between the learned feature spaces of different experts, which is also potentially reflected by their weights.
Based on this analysis, we can use the weights as the expert representation to discover similar experts for pruning. In our implementation, we consider two strategies: (1) directly using the vectorized weights, which does not account for the structure of experts, and (2) leveraging the local linearity of neural networks to compute the approximated weight matrix, namely the surrogate weights in our paper.
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$.
Q3: How do we determine the value of $\alpha$ for merging similar experts?
Two strategies are explored in our implementation to merge similar experts.
One strategy is to simply merge experts and gates with the same coefficients, i.e., $\alpha_{1}=\alpha_{2}=...=\alpha_{|\mathcal{V}_{n}|}=\frac{1}{|\mathcal{V}_{n}|}$, which encourages experts contributing equally.
The other strategy is to learn the merging coefficients for different experts towards minimizing the following loss function,
($\mathcal{L}(\{\alpha_{n}\}_{n=1}^{|\mathcal{V}_{n}|})=\|y-F(x;\hat{\Theta},\hat{W},K)\|,$ (7)
[…]
$\hat{W}_{n}=\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups.
4 Evaluation
Table 1: Results on pruning the Mixtral-8x7B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Augmented: prunable experts discover using augmented data and merging with uniform coefficients; 3) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Router-guided [21] 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
Count-guided [11] 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
Enumerate [27] 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
Ours Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
Augmented 56.2/47.8 72.7/57.2 52.1/42.7 68.1/54.0 85.6/83.2 32.8/28.6 68.6/66.5 62.3/54.3
Learn 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
4.1 Experiment Setup
Studied Models. We take pruning Mixtral-8x7B and Mixtral-8x22B [14] as the example. The Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The hidden dimension for token embedding is $4,096$, and the intermediate feature dimension in each FFN of MoEs is $14,336$. The Mixtral-8x22B is similar with Mixtral-8x7B, but with $56$ sparse MoE layers. It has the token embedding dimension of $6,144$ and the intermediate feature dimension of $16,384$. During the inference, each token will select $2$ experts in each MoE layer.
Compared Baselines. We take three advanced MoE pruning methods [11, 21, 27] as our baseline for comparison. Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [27], also the count-guided strategy and loss-based pruning method [27], namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
Our Approaches. We use four metrics to prune the model, including 1) Vectorized $\theta$: concact the vectorized model parameter $\theta$ together as the expert representation for grouping, and apply the uniform merging on the experts and routers. 2) Surrogate $\theta$: multiply the model parameters together as the weight for one-layer surrogate FFN and use it as the expert representation. The merging strategy is uniform merging. 3) Augmented: compute the expert representation based on the expert outputs on the augmented inputs. The merging strategy is uniform merging. 4) Learn: Use the same grouping strategy with Augmented but learn to merge for experts and routers.
Evaluation Datasets. The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance. We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [31], and RTE [1]. Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
Implementations. We randomly select $128$ samples from the pre-training dataset C4 [32] as the calibration dataset. For the prunable candidate group discovery, we use the mean square error to evaluate the similarity between experts. For the learnable merging, we initialize the merging coefficients $\alpha$ as the uniform and set $\lambda$ as $1$. In our experiments, we utilize four NVIDIA A100 GPUs, each with 80GB of memory.
4.2 Pruning the Mixtral Architecture
We present the results of pruning the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B, from 8 experts to 6 and 4 experts in each layer, respectively. The results on the two models are depicted in table 1 and table 2 respectively.
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies [27] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning. 2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance. Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
Results on Mixtral-8x22B. With more parameters introduced in Mixtral-8x22B, the search for optimal experts to drop and merge becomes more challenging. Unlike the results on pruning Mixtral-8x7B, the strategy using the surrogate $\theta$ achieves state-of-the-art results, with only a minor performance drop of $2.8\%$ on average compared to the full model.
At the same time, it should be noted that the use of data doesn’t help boost the pruning performance as well, where the augmented and learning-based method has an average performance drop of $1.05\%$ and $1.0\%$, respectively. Also, the routing similarity-based pruning suffers model collapse under both two settings. We argue it is due to the heavy imbalance between the used data and parameter space, leading to the computation of expert similarity over-fitted on the small calibration dataset.
Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$, 2) Augmented, and 3) Learn. The first column indicate the results on pruned model with $6$ experts, while the second column present the result of $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Router-guided [21] 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
Count-guided [11] 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
Enumerate [27] 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
Ours Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
Augmented 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
Learn 61.4/57.9 78.3/70.1 61.2/51.4 72.8/65.0 88.2/84.9 35.6/32.7 70.5/67.3 66.9/61.3
4.3 Empirical Analysis on Expert Hints
While expert pruning brings a significant speedup by reducing the parameters, what changes are introduced to different MoE layers? To answer this question, we analyze the change of expert hints on the calibration dataset C4.
Figure 4: Statistics of the visiting frequency for all experts in different MoE layers.
The statistic results are shown in fig. 4. We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints. In contrast, pruning differentiates the hints of experts. Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints. Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
4.4 Ablation Study and Discussion
On the number of samples. In our work, we augment the data to improve the accuracy of similarity estimation. While this technique is typically used in scenarios with limited data, a natural question arises:
Would increasing the amount of data help preserve performance after pruning?
Table 3: Evaluation results of different pruning setting with ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
# of samples MMLU BoolQ OpenBookQA RTE Avg.
128 0.615 0.856 0.328 0.686 0.621
256 0.614 0.854 0.332 0.697 0.624
512 0.617 0.854 0.334 0.704 0.627
2,048 0.612 0.856 0.322 0.675 0.616
We conduct experiments with different numbers of samples to answer this question. As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$. It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
We attribute this phenomenon to the use of mean operation in our expert similarity estimation. On the one hand, the mean operation keeps the feature dimension for efficient pairwise distance computation. On the other hand, it naturally downgrades the precision of function similarity computation by “round-off” error compared with other per sample-based distance metrics like CKA used in section 3.2.
On the choice of distance function. We also considered other distance functions, such as cosine similarity, to evaluate expert similarity during the discovery stage. However, it performed worse compared to our mean-squared-error function, showing a drop of $0.9\%$ on MMLU, $3.3\%$ on RTE, and $0.2\%$ on BoolQ.
On merging the routing policy. One relevant work to ours is [21], which identifies and merges similar experts based on routing policy, setting the routing policy of the merged expert to that of the most frequently visited expert. While we identify similar experts based on expert knowledge, such as weights and features, we merge not only the experts but also the routing policies. To demonstrate the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies. As shown in table 4, this results in performance degradation across all tasks and different numbers of experts, highlighting the crucial role of simultaneously merging the routers and experts.
Table 4: Evaluation results of our augmentation-based pruning method without merging the routers.
# of experts MMLU BoolQ OpenBookQA RTE Avg.
4 0.491 0.833 0.274 0.664 0.565
6 0.604 0.847 0.312 0.675 0.609
5 Conclusion and Future Work
In this paper, we address the practical problem of reducing the number of experts in MoE layers to accelerate the inference and fine-tuning process. We propose discovering similar experts at the feature level and then merging them in the weight space for MoE pruning while preserving as much original expert knowledge as possible. This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating significant potential for further pruning. Future work will aim to push the compression of MoE layers to the limit.
6 Acknowledgments
We would like to thank Lifeng Li from Microsoft for setting up GPU machines.
Appendix A Evaluation on the expert similarity
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. A1 and table 2. We summarize the observation as follows,
  • Most MoE layers in the two Mixtral models contain significant expert redundancy.
  • The most redundant MoE layers are located in the first and last several layers, while the experts in the intermediate MoE layers learn more diverse features.
Figure A1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
Figure A2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.

### v2

@@
 By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost.
-However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real world settings.
+However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real-world settings.
 Our empirical study reveals that some experts encode redundant knowledge during pre-training.
-We thus propose a method of grouping and pruning similar experts to improve model’s parameter efficiency.
-We validate the effectiveness of our method by pruning two state-of-the-art MoE models, Mixtral-8x7B and Mixtral-8x22B.
-Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
-To facilitate future research, we will release our code and the pruned MoE models.
+We thus propose a method of grouping and pruning similar experts to improve the model’s parameter efficiency.
+We validate the effectiveness of our method by pruning three state-of-the-art MoE architectures, including Mixtral, Deepseek-MoE, and Qwen.
+The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
+We will release our code to facilitate future research.
 […]
@@
 In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
-There are many recent studies on pruning MoE models.
+There have been several studies on pruning MoE models.
 One line of work utilizes task-specific information to prune irrelevant experts.
@@
 Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers.
-We validate the effectiveness of our method by pruning experts for two state of the art MoE models, Mixtral-8x7B and Mixtral-8x22B.
-Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
+We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A).
+The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
 2 Related Work
@@
 3.1 Task Definition
-Several prior works [27, 11, 21] have identified that removing some experts in different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks.
+Previous works [26, 11, 21] have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks.
 The remaining question is which expert(s) should be pruned in the MoE layer.
-Specifically, for a MoE LLM $\mathcal{M}(\cdot;\mathcal{F})$, which consists of multiple MoE layers =$\mathcal{F}=\{F^{l}(\cdot;\Theta^{l},W^{l},K)\}_{l=1}^{L}$, we want to prune $r$ experts from the total $N$.
+Specifically, for the $l$-th MoE layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ in the given LLM $\mathcal{M}(\cdot;\mathcal{F})$, we want to prune $r$ experts from the total of $N$, $l=1,2,...,L$.
 This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete.
-We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ that satisfy the following optimization objective:
+We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ towards the following optimization objective:
 $\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y),$ (3)
-where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the pruned LLM by deleting $r$ experts from each MoE layer $F$ in $\mathcal{F}$, and $\mathcal{D}$ is the dataset.
-This task is particularly challenging as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions.
-In this work, we propose to estimate expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
+where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the new LLM with pruned MoE layer $\hat{\mathcal{F}}$, and $\mathcal{D}$ is any dataset under the task-agnostic setting.
+This task is particularly challenging, as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions.
+In this work, we propose to estimate the expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
 Figure 2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under two kernel-based CKA criteria (Linear and RBF).
@@
 For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
-Remark.
-We notice that a concurrent work [26] also studies expert similarity, but focuses on the cosine similarity between expert weights.
-Our work examines the differences between experts using a more generalized metric based on the correlation between inputs and outputs, and obtains a different observation: even without considering the last layer, expert similarity does not gradually decrease with increasing depth of layers.
-For example, we observe higher pairwise expert similarity at the $11$-th layer compared to the $6$-th layer, which is cross-verified by both distance metrics.
-Additionally, the experts in the $26$-th layer also demonstrate a higher similarity with others compared with shallow layers.
-We provide the full evaluation results on visualizing the expert similarity of MoE layers in Mixtral-8x7B and Mixtral-8x22B in appendix A.
 3.3 Discover and Merge Similar Experts on Feature Space
@@
 Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
-Then, we initialize the $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
-Last, we leverage the graph partition algorithm to split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$.
-The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ shares the most similar knowledge with each other, and shows much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$.
-To achieve this goal, we adopt a two-step strategy.
-At the first step, we partition the $\mathcal{V}$ of $\mathcal{G}$ into two subsets $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, where $\mathcal{V}_{1}$ has $r-1$ nodes and $\mathcal{V}_{2}$ has $G-r+1$ nodes.
-Then, we split $\mathcal{V}_{1}$ into $r-1$ subsets, $\mathcal{V}_{1,1},...,\mathcal{V}_{1,r-1}$.
-The splitting process can be formulated as follows,
-$\min\sum_{i,j\in\mathcal{V}}x_{i}x_{j}D_{ij}-\sum_{i,j\in\mathcal{V}}x_{i}(1-x_{j})D_{ij}$ (5)
-$-\sum_{i,j\in\mathcal{V}}(1-x_{i})(1-x_{j})D_{ij},$
+Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
+Last, we split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$ to group similar experts.
+The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ share the most similar knowledge with each other and show much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$.
+This can be formulated as the follows,
+$\min\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}D_{jk}-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}D_{jk}\right),$ (5)
 […]
-where $x_{i}=1$ if node $i$ in group $\mathcal{V}_{1}$, and $x_{i}=0$ otherwise.
-The first term is used to minimize the distance between nodes in $\mathcal{V}_{1}$, the second term is usd to maximize the distance of nodes between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, and the third term is ued to maximize the distance of nodes in $\mathcal{V}_{2}$.
-When pruning $G$ experts to $r$ experts, there are $r-1$ experts keeping unaffected in the entire process.
-Our graph partition strategy can preserve the most knowledge while diversifying the expert features in each MoE layers.
+where The objective function seeks to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $D_{jk}$ based on expert $i$ and $j$.
 Merging Similar Experts with Routers.
-To preserve the different knowledge learned by different models clustered in the same group, we merge these experts as well as their routers on the weight space as follows,
-$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},$ (6)
-$\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
+To diversity the experts and preserve the different knowledge learned by different models clustered in the same group [37, 35], we merge the clusted experts with their routers on the weight space as follows,
+$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},~{}~{}\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$ (6)
 where $\mathcal{V}_{n}$ is the set of similar experts in the $n$-th cluster $\mathcal{G}_{n}$, and we have $\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}=1$.
@@
 3.4 Practical Implementation
-While we have developed a general strategy to prune sparse MoEs, several practical questions arise during implementation:
-Q1: How to decide $\mathcal{M}(\cdot)$ and perform task-agnostic pruning on sparse MoEs?
-Intuitively, given the same input, similar experts should generate similar outputs.
-Thus, we propose to use the mean squared error on the outputs of two experts to evaluate the similarity between them.
-We follow previous work [27] to select few samples from pre-trained dataset C4 [32] as the calibration dataset to compute the expert feature representation at the third step for similar experts discovery.
-The C4 dataset comprises multiple domains of contents, which can be a good surrogate to simulate the performance of the task-agnostic pruning.
-However, with the limited size of the calibration dataset, the estimation of expert similarity based on the aforementioned expert representation could be inaccurate due to biased sampling.
-To improve the generalization ability at the discovery stage, we randomly mix up the token embedding for data augmentation and then use the augmented data for computing the expert representation and similarity.
-In practice, we iteratively update the expert similarity matrix by continuously shuffling and mixing up the data for full augmentation.
-Q2: Is it possible to prune sparse MoEs without using any data?
-Experts in the MoE layer of current models typically have the same structure; for instance, in Mixtral-8x7B, each expert is a two-layer FFN with SiLU activation.
-By training on different routed tokens, the experts usually learn different distributions during the training process.
-This results in a disparity between the learned feature spaces of different experts, which is also potentially reflected by their weights.
-Based on this analysis, we can use the weights as the expert representation to discover similar experts for pruning.
-In our implementation, we consider two strategies: (1) directly using the vectorized weights, which does not account for the structure of experts, and (2) leveraging the local linearity of neural networks to compute the approximated weight matrix, namely the surrogate weights in our paper.
+On the discovery step. We propose two strategies for computing the disparity matrix $D$, including the data-centric strategy and model-centric strategy.
+For data-centric strategy, we follow previous work [26] to use the pre-trained dataset C4, which can be a good surrogate to simulate the performance of the task-agnostic pruning.
+To improve the generalization ability at the discovery stage, we randomly mix the token embedding for data augmentation and then use the augmented data to calculate the expert output as the expert representation.
+For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation.
+The first way is to leverage the vectorized weights directly.
+The other way is to leverage the local linearity of neural networks to compute the surrogate weight matrix.
 Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$.
@@
 Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$.
-Q3: How do we determine the value of $\alpha$ for merging similar experts?
-Two strategies are explored in our implementation to merge similar experts.
-One strategy is to simply merge experts and gates with the same coefficients, i.e., $\alpha_{1}=\alpha_{2}=...=\alpha_{|\mathcal{V}_{n}|}=\frac{1}{|\mathcal{V}_{n}|}$, which encourages experts contributing equally.
-The other strategy is to learn the merging coefficients for different experts towards minimizing the following loss function,
+Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance in pruning models (see section A.2).
+On the merging step.
+We consider three approaches in practice deciding $\alpha$ to merge similar grouped experts.
+The first is to only maintain the experts with the maximum visiting frequency and drop all the others.
+The second is to set $\alpha_{i}=\frac{1}{|\mathcal{V}_{n}|}$, which uniformly assembles all the experts grouped.
+The last one is learning $\alpha$ to merge the grouped experts by minimizing the following loss function,
 ($\mathcal{L}(\{\alpha_{n}\}_{n=1}^{|\mathcal{V}_{n}|})=\|y-F(x;\hat{\Theta},\hat{W},K)\|,$ (7)
 […]
-$\hat{W}_{n}=\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
 where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups.
+Among these three strategies, we empirically find that the uniform souping strategy offers more stable performance and greater efficiency in model pruning (see section A.2 ).
+While the learning strategy can yield slightly better performance, it is more time-consuming due to the need for tuning the parameter $\alpha$.
 4 Evaluation
 Table 1: Results on pruning the Mixtral-8x7B from 8 experts to 6 and 4 experts in each MoE layer.
-We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Augmented: prunable experts discover using augmented data and merging with uniform coefficients; 3) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq.
-7.
-The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
+The first and second columns respectively indicate the results of the pruned model with $6$ and $4$ experts.
 Dataset MMLU BoolQ OpenBookQA RTE Average
@@
 Enumerate [26] 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
-Ours Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
-Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
-Augmented 56.2/47.8 72.7/57.2 52.1/42.7 68.1/54.0 85.6/83.2 32.8/28.6 68.6/66.5 62.3/54.3
-Learn 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
+Ours Model-centric 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
+Data-centric 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
 4.1 Experiment Setup
 Studied Models.
-We take pruning Mixtral-8x7B and Mixtral-8x22B [14] as the example.
-The Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts.
-The hidden dimension for token embedding is $4,096$, and the intermediate feature dimension in each FFN of MoEs is $14,336$.
-The Mixtral-8x22B is similar with Mixtral-8x7B, but with $56$ sparse MoE layers.
-It has the token embedding dimension of $6,144$ and the intermediate feature dimension of $16,384$.
+We take pruning three MoE-based architectures as an example, including the Mixtral, Deepseek, and Qwen.
+For the Mixtral architecture, the Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts.
+The Mixtral-8x22B is similar to Mixtral-8x7B but with $56$ sparse MoE layers.
 During the inference, each token will select $2$ experts in each MoE layer.
-Compared Baselines.
+The deepseek model has $28$ layers, and there are $64$ experts in each layer.
+Each token will pass $2$ shared experts and select $6$ experts during the inference.
+The Qwen model also has $28$ MoE layers with $64$ experts in each layer but will activate $8$ experts during the inference.
+Pruning Methods.
 We take three advanced MoE pruning methods [11, 21, 26] as our baseline for comparison.
@@
 Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
-Our Approaches.
-We use four metrics to prune the model, including 1) Vectorized $\theta$: concact the vectorized model parameter $\theta$ together as the expert representation for grouping, and apply the uniform merging on the experts and routers.
-2) Surrogate $\theta$: multiply the model parameters together as the weight for one-layer surrogate FFN and use it as the expert representation.
-The merging strategy is uniform merging.
-3) Augmented: compute the expert representation based on the expert outputs on the augmented inputs.
-The merging strategy is uniform merging.
-4) Learn: Use the same grouping strategy with Augmented but learn to merge for experts and routers.
-Evaluation Datasets.
-The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance.
-We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [31], and RTE [1].
-Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
-Implementations.
-We randomly select $128$ samples from the pre-training dataset C4 [32] as the calibration dataset.
-For the prunable candidate group discovery, we use the mean square error to evaluate the similarity between experts.
-For the learnable merging, we initialize the merging coefficients $\alpha$ as the uniform and set $\lambda$ as $1$.
-In our experiments, we utilize four NVIDIA A100 GPUs, each with 80GB of memory.
-4.2 Pruning the Mixtral Architecture
-We present the results of pruning the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B, from 8 experts to 6 and 4 experts in each layer, respectively.
-The results on the two models are depicted in table 1 and table 2 respectively.
-Results on Mixtral-8x7B.
-We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average.
-Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge.
-Besides, compared with count-guided and enumerate strategies [27] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
-From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods.
-This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
-2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance.
-Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
-Results on Mixtral-8x22B.
-With more parameters introduced in Mixtral-8x22B, the search for optimal experts to drop and merge becomes more challenging.
-Unlike the results on pruning Mixtral-8x7B, the strategy using the surrogate $\theta$ achieves state-of-the-art results, with only a minor performance drop of $2.8\%$ on average compared to the full model.
-At the same time, it should be noted that the use of data doesn’t help boost the pruning performance as well, where the augmented and learning-based method has an average performance drop of $1.05\%$ and $1.0\%$, respectively.
-Also, the routing similarity-based pruning suffers model collapse under both two settings.
-We argue it is due to the heavy imbalance between the used data and parameter space, leading to the computation of expert similarity over-fitted on the small calibration dataset.
+For our method, we respectively report the best results of our data-centric and model-centric methods in pruning models.
+Following the previous setting [26], we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data.
+More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in section A.2.
 Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer.
-We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$, 2) Augmented, and 3) Learn.
-The first column indicate the results on pruned model with $6$ experts, while the second column present the result of $4$ experts.
+The first and second columns respectively indicate the results of the pruned model with $6$ and $4$ experts.
 Dataset MMLU BoolQ OpenBookQA RTE Average
@@
 Enumerate [26] 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
-Ours Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
-Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
-Augmented 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
-Learn 61.4/57.9 78.3/70.1 61.2/51.4 72.8/65.0 88.2/84.9 35.6/32.7 70.5/67.3 66.9/61.3
+Ours Model-centric 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
+Data-centric 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
+Evaluation Datasets.
+The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance.
+We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [30], and RTE [1].
+Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
+4.2 Pruning the Mixtral Architecture
+We present the results of the pruning of the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B.
+Both have $8$ experts in each MoE layer.
+We apply different methods to prune them from 8 experts to 6 and 4 experts in each layer, respectively.
+The results of the two models are shown in table 1 and table 2 respectively.
+Results on Mixtral-8x7B.
+We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average.
+Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge.
+Besides, compared with count-guided and enumerate strategies [26] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
+We can also notice that the model-centric method surpasses all the other data-involved pruning baseline methods.
+This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
+Results on Mixtral-8x22B.
+In this experiment, the model-centric method achieves the best result with only a minor performance drop of $2.8\%$ on average compared to the full model.
+Our proposed data-centric method achieves the second to the last performance, with a gap of $0.8\%$ performance gap compared with the runner-up method.
+It should be noted that the routing similarity-based pruning suffers model collapse under both two settings, showing a less robustness compared with other methods and metrics for similarity computation.
+4.3 Pruning the Deepseek Model
+We also evaluate the effectiveness of our proposed pruning method on compressing the Deepseek architecture.
+Specifically, we pruned the experts of Deepseek-MoE-16B, reducing the number of experts from 64 to 48 using different model pruning strategies.
+The results are illustrated in fig.
+4.
+Notably, even after pruning one-third of the experts in Deepseek-MoE-16B, our data-centric strategy maintains an impressive average performance of $50.9\%$, with only a $3.1\%$ reduction in performance compared to the full model.
+In the evaluation of the most challenging MMLU task, our model-centric strategy demonstrates superior performance in most cases, particularly when reducing the number of experts from $62$ to $52$.
+It consistently outperforms the runner-up baseline method, achieving a clear performance advantage of $2.6\%$.
+Figure 4: Results on pruning the Deepseek-MoE-16B with different strategies.
+4.4 Pruning the Qwen Model
+Figure 5: Results on pruning the Qwen2-57B-14A with different strategies.
+The Qwen architecture is also utilized in our experiments.
+We study the Qwen2-57B-14A to evaluate the performance of different model pruning strategies.
+We prune the experts from $64$ experts to $48$ experts in each layer and evaluate the performance on different tasks.
+The results are reported in fig.
+5.
+When pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method.
+On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the count-guided pruning strategy.
 4.5 Empirical Analysis on Expert Hints
@@
 The statistic results are shown in fig.
-4.
+6.
 We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints.
@@
 Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
-4.4 Ablation Study and Discussion
-On the number of samples.
-In our work, we augment the data to improve the accuracy of similarity estimation.
-While this technique is typically used in scenarios with limited data, a natural question arises:
-Would increasing the amount of data help preserve performance after pruning?
 Table 3: Evaluation results of different pruning setting with ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
@@
 2,048 0.612 0.856 0.322 0.675 0.616
-We conduct experiments with different numbers of samples to answer this question.
+4.6 Ablation Study and Discussion
+Comparison with the greedy search.
+To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method.
+Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer.
+We then merge these results to prune the model across all layers.
+The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result.
+We provide the full enumeration results in section A.1.
+On the number of samples.
+We conduct experiments to study the effect of the used sample in data-centric pruning methods.
 As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$.
@@
 On the other hand, it naturally downgrades the precision of function similarity computation by “round-off” error compared with other per sample-based distance metrics like CKA used in section 3.2.
-On the choice of distance function.
-We also considered other distance functions, such as cosine similarity, to evaluate expert similarity during the discovery stage.
-However, it performed worse compared to our mean-squared-error function, showing a drop of $0.9\%$ on MMLU, $3.3\%$ on RTE, and $0.2\%$ on BoolQ.
 On merging the routing policy.
@@
 However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers.
-Our analysis reveals that some experts have low visitation frequencies, indicating significant potential for further pruning.
-Future work will aim to push the compression of MoE layers to the limit.
+Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
 6 Acknowledgments
@@
 The results are respectively depicted in fig.
-A1 and table 2.
+A1 and fig.
+A2.
 We summarize the observation as follows,
@@
 Figure A2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.
+A.1 Enumeration on the expert pruning
+We present the full greedy search result on Mixtral-8x7B in fig.
+A3.
+In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task.
+For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
+Although dropping most of the combinations on different layers only leads to a minor performance drop, we can notice that it could cause the model to crash when the fourth expert in layer 1 is involved during the pruning process.
+Figure A3: Enumeration on dropping two experts.
+A.2 Results on different strategies for pruning
+While we report the best performance using our data-centric and model-centric strategies to prune the MoE models, we detail more results on pruning the Mixtral-8x7B and Mixtral-8x22B.
+For model centric, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts.
+We uniformly merge the experts and their routers for model pruning.
+For data centric, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), or only maintain the expert with the maximum visiting frequency (Max).
+From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods.
+This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
+2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance.
+Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
+Table A1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer.
+We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq.
+7; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency.
+The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
+Dataset MMLU BoolQ OpenBookQA RTE Average
+Method humanities social science stem Other
+Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
+Model-centric Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
+Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
+[…]
+Max 56.4/47.6 71.9/58.6 52.0/42.9 66.9/55.7 85.1/82.8 35.2/28.8 70.4/66.8 62.6/53.7
+Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
+Model-centric Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
+Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
+[…]
+Max $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7

### v3

@@
 1 Introduction
-Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws [16].
+Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws (Kaplan et al., 2020).
 In addition to increasing the depth of neural network models, widening neural networks by using the sparsely-activated mixture-of-experts (MoE) architecture is also proved effective.
@@
 Thus, compared to dense models, MoE models achieve better end-task performance and generalize better to new tasks without increasing computation costs.
-Notable examples of MoE models include Switch Transformer [9], Mixtral-MoE [14], and Uni-MoE [22].
+Notable examples of MoE models include Switch Transformer (Fedus et al., 2022), Mixtral-MoE (Jiang et al., 2024), and Uni-MoE (Li et al., 2024b).
 Despite significant progress in developing wider and deeper MoE LLMs, the increased memory consumption due to larger model sizes (i.e., increased number of experts) poses a substantial challenge to the deployment of these models in real-world settings.
@@
 Adding or removing even one expert in each layer can significantly impact overall memory cost and model performance.
-For example,  [26] shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark [12].
+For example,  (Lu et al., 2024) shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark (Hendrycks et al., 2020).
 In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
@@
 For example, Chen et al.
-[3] prune the less frequently visited experts based on experiments on a range of tasks.
+(2022) prune the less frequently visited experts based on experiments on a range of tasks.
 Chowdhury et al.
-[5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage.
+(2024) find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage.
 Li et al.
-[21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning.
+(2024a) merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning.
 Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks.
 He et al.
-[11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop.
+(2024) explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop.
 In comparison, Lu et al.
-[26] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance.
+(2024) enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance.
 We illustrate the difference among these works in fig.
@@
 Figure 1: Removing several experts from the original MoE layer would not cause model collapse but improve efficiency.
+Experts with similar colors share the similar knowledge with each other.
 Prior works often utilize expert access information to filter out unimportant experts.
@@
 As shown in fig.
-1, we first identify and group similar experts in the feature space.
+3, we first identify and group similar experts in the feature space.
 Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers.
@@
 The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
+Our contribution can be summarized as follows,
+  1. 1.
+    We empirically validate that some experts within each well-trained MoE layer encode similar knowledge, making them somewhat redundant.
+  2. 2.
+    We propose a two-stage, task-agnostic method for grouping and merging redundant experts, which is further divided into data-centric and model-centric implementation strategies.
+  3. 3.
+    We demonstrate the effectiveness of our method by pruning experts from a series of state-of-the-art MoE models, including Mixtral-MoE, DeepSeek-MoE, and Qwen-MoE.
+The results from a greedy search for MoE pruning further validate the success of our approach.
 2 Related Work
@@
 The Mixture-of-Experts (MoE) structure is firstly applied in classical machine learning models by Jacobs et al.
-[13] and Jordan and Jacobs [15], then widely used in various deep learning models [39, 29, 40].
-Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models [32, 19, 43].
+(1991) and Jordan and Jacobs (1994), then widely used in various deep learning models (Yuksel et al., 2012; Masoudnia and Ebrahimpour, 2014; Zhang et al., 2023).
+Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models (Shazeer et al., 2017; Lepikhin et al., 2020; Zoph et al., 2022).
 It adapts the original large feed-forward network (FFN) in each transformer block into multiple smaller FFNs, forming an expert layer with a router that computes the weighted output of each MoE layer.
 Sparse MoEs were first proposed by Fedus et al.
-[9].
+(2022).
 In this approach, only a few experts are activated in each layer, accelerating training and inference while significantly increasing the number of parameters for greater model capacity.
-Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer [9], Mixtral-8B [14], and Uni-MoE [22].
-Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models [41, 20].
+Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer (Fedus et al., 2022), Mixtral-8B (Jiang et al., 2024), and Uni-MoE (Li et al., 2024b).
+Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models (Zhu et al., 2022; Li et al., 2022).
 Model Pruning.
-Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency [25, 36, 24].
-Pruning techniques can be categorized into unstructured pruning [23, 34, 28], which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning [18, 8, 33], which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs [4].
-Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs [27], RNNs [42], and LSTMs [7] to modern large models such as Llama [38] and Stable-Diffusion [2].
+Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency (Liu et al., 2018; Wang et al., 2020; Liu et al., 2021).
+Pruning techniques can be categorized into unstructured pruning (Liao et al., 2023; Shi et al., 2024; Mason-Williams and Dahlqvist, 2024), which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning (Lemaire et al., 2019; Fang et al., 2023; Shen et al., 2022), which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs (Choi and Yang, 2021).
+Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs (Luo et al., 2018), RNNs (Zhu and Gupta, 2017), and LSTMs (Ding et al., 2020) to modern large models such as Llama (Xia et al., 2023) and Stable-Diffusion (Castells et al., 2024).
 While the large amount of parameters in sparse MoEs benefits the model’s capacity to achieve good performance at the pre-training stage, the increasing memory consumption causes great challenges to fine-tuning different downstream tasks.
@@
 3.1 Task Definition
-Previous works [26, 11, 21] have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks.
+Previous works (Lu et al., 2024; He et al., 2024; Li et al., 2024a) have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks.
 The remaining question is which expert(s) should be pruned in the MoE layer.
@@
 Conversely, the redundancy of shared knowledge learned by different experts provides opportunities for model pruning.
-We employ the Centered Kernel Alignment (CKA) [17] as the criteria to evaluate the similarity between experts in each MoE layer.
+We employ the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) as the criteria to evaluate the similarity between experts in each MoE layer.
 Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
@@
 where $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
-Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 [31].
+Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 (Raffel et al., 2020).
 We respectively use the linear kernel and RBF kernel in similarity computation of eq.
@@
 For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
+Full evaluation results can be found in appendix B.
 3.3 Discover and Merge Similar Experts on Feature Space
-Figure 3: Overview of our method.
-We first leverage different metrics to obtain the expert representation, then compute the pairwise distance to get the expert similarity matrix.
-Based on the expert similarity matrix, we can group the experts with shared knowledge in the same cluster, which can be merged on the weight space for MoE pruning.
+Figure 3: We first leverage model or data-centric strategies to obtain the expert representation, then compute the pairwise distance to get the disparity matrix.
+Based on the expert similarity matrix, we can group similar experts with shared knowledge in the same cluster, which can be merged on the weight space for pruning.
 Having identified the existence of moderate expert similarity in each MoE layer, we propose pruning the experts within each MoE layer by merging experts with shared knowledge.
@@
 Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
-Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
+Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the pair-wise expert distance matrix $D$.
 Last, we split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$ to group similar experts.
@@
 Merging Similar Experts with Routers.
-To diversity the experts and preserve the different knowledge learned by different models clustered in the same group [37, 35], we merge the clusted experts with their routers on the weight space as follows,
+To diversity the experts and preserve the different knowledge learned by different models clustered in the same group (Wortsman et al., 2022; Wang et al., 2022), we merge the clusted experts with their routers on the weight space as follows,
 $\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},~{}~{}\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$ (6)
@@
 On the discovery step. We propose two strategies for computing the disparity matrix $D$, including the data-centric strategy and model-centric strategy.
-For data-centric strategy, we follow previous work [26] to use the pre-trained dataset C4, which can be a good surrogate to simulate the performance of the task-agnostic pruning.
-To improve the generalization ability at the discovery stage, we randomly mix the token embedding for data augmentation and then use the augmented data to calculate the expert output as the expert representation.
+For the data-centric strategy, since the full pre-training datasets for different models are large and inaccessible, we use the C4 dataset, a commonly used smaller pre-training subset that serves as an effective surrogate for capturing task-agnostic knowledge of experts Lu et al.
+(2024).
+Specifically, by disabling the router function, for the same input, we use the output of each expert as its expert representation $\mathcal{M}(\cdot)$.
+To enhance generalization and mitigate overfitting to the selected samples during model pruning, we apply data augmentation by randomly mixing token embeddings during the representation computation at the discovery stage.
 For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation.
-The first way is to leverage the vectorized weights directly.
-The other way is to leverage the local linearity of neural networks to compute the surrogate weight matrix.
+One is to leverage the vectorized weights directly ($\clubsuit$).
+The other is to leverage the local linearity of neural networks (Zhang and Wu, 2020) to compute the surrogate weight matrix ($\spadesuit$).
 Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$.
 For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$.
-Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$.
-Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance in pruning models (see section A.2).
+Then, the vectorized weight ($\clubsuit$) is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$ ($\spadesuit$).
+Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance (see appendix D).
 On the merging step.
@@
 where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups.
-Among these three strategies, we empirically find that the uniform souping strategy offers more stable performance and greater efficiency in model pruning (see section A.2 ).
+Among these three strategies, both the first and third require the presence of data, while the second is compatible with both data-centric and model-centric approaches.
+We empirically find that the uniform souping strategy offers more stable performance and greater efficiency in task-agnostic model pruning (see appendix D ).
 While the learning strategy can yield slightly better performance, it is more time-consuming due to the need for tuning the parameter $\alpha$.
@@
 Method humanities social science stem Other
-Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
-Router-guided [21] 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
-Count-guided [11] 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
-Enumerate [26] 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
+Mixtral-8x7B (Jiang et al., 2024) 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
+Router-guided (Li et al., 2024a) 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
+Count-guided (He et al., 2024) 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
+Enumerate (Lu et al., 2024) 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
 Ours Model-centric 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
@@
 Pruning Methods.
-We take three advanced MoE pruning methods [11, 21, 26] as our baseline for comparison.
-Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset.
-For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [26], also the count-guided strategy and loss-based pruning method [26], namely the Enumerate in our paper.
+We take three advanced MoE pruning methods (He et al., 2024; Li et al., 2024a; Lu et al., 2024) as our baseline for comparison.
+Among them, router-guided merging (Li et al., 2024a) is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset.
+For task-agnostic methods, we select the frequency-based pruning method Expert Trimming (Lu et al., 2024), also the count-guided strategy and loss-based pruning method (Lu et al., 2024), namely the Enumerate in our paper.
 Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
 For our method, we respectively report the best results of our data-centric and model-centric methods in pruning models.
-Following the previous setting [26], we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data.
-More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in section A.2.
+Following the previous setting Lu et al.
+(2024), we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data.
+More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in appendix D.
 Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer.
@@
 Method humanities social science stem Other
-Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
-Router-guided [21] 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
-Count-guided [11] 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
-Enumerate [26] 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
+Mixtral-8x22B (Jiang et al., 2024) 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
+Router-guided (Li et al., 2024a) 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
+Count-guided (He et al., 2024) 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
+Enumerate (Lu et al., 2024) 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
 Ours Model-centric 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
@@
 Evaluation Datasets.
-The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance.
-We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [30], and RTE [1].
-Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
+The open-sourced Language Model Evaluation Harness library (Gao et al., 2021) is used to evaluate the performance.
+We select four tasks, including MMLU (Hendrycks et al., 2020), BoolQ (Clark et al., 2019), OpenBookQA (Mihaylov et al., 2018), and RTE (Bentivogli et al., 2009).
+Among these tasks, MMLU is the most challenging one, which consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
 4.2 Pruning the Mixtral Architecture
@@
 We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average.
-Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge.
-Besides, compared with count-guided and enumerate strategies [26] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
+Compared with count-guided strategy He et al.
+(2024) which just drops the experts less visited, router-guided strategy (Li et al., 2024a) has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge.
+Besides, compared with count-guided and enumerate strategies (Lu et al., 2024) which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
 We can also notice that the model-centric method surpasses all the other data-involved pruning baseline methods.
@@
 Results on Mixtral-8x22B.
-In this experiment, the model-centric method achieves the best result with only a minor performance drop of $2.8\%$ on average compared to the full model.
-Our proposed data-centric method achieves the second to the last performance, with a gap of $0.8\%$ performance gap compared with the runner-up method.
-It should be noted that the routing similarity-based pruning suffers model collapse under both two settings, showing a less robustness compared with other methods and metrics for similarity computation.
+In this experiment, the model-centric approach achieves the best result, with only a minor performance drop of $2.8\%$ on average compared to the full model.
+Our proposed data-centric method ranks second to last, with a performance gap of $0.8\%$ compared to the runner-up method.
+This suggests that model-centric approaches exhibit better robustness when pruning models of different scales, while data-centric methods are more prone to overfitting on small calibration datasets (as evidenced by the collapse of the route-guided method in table 2).
 4.3 Pruning the Deepseek Model
+Figure 4: Results on pruning the Deepseek-MoE-16B with different strategies.
 We also evaluate the effectiveness of our proposed pruning method on compressing the DeepSeek architecture.
-Specifically, we pruned the experts of Deepseek-MoE-16B, reducing the number of experts from 64 to 48 using different model pruning strategies.
+Unlike Mixtral-MoE, DeepSeek-MoE features a shared expert and incorporates more fine-grained experts at each MoE layer.
+Specifically, we pruned the non-shared experts of DeepSeek-MoE-16B, reducing the number of experts from 64 to 48 using various model pruning strategies.
 The results are illustrated in fig.
-4.
+4***We only show the pruning results of the Deepseek and Qwen models on the most challenging MMLU task and the average performance across MMLU, BoolQ, OpenBookQA, and RTE.
+Full results can be found in supplementary..
 Notably, even after pruning one-third of the experts in Deepseek-MoE-16B, our data-centric strategy maintains an impressive average performance of $50.9\%$, with only a $3.1\%$ reduction in performance compared to the full model.
@@
 It consistently outperforms the runner-up baseline method, achieving a clear performance advantage of $2.6\%$.
-Figure 4: Results on pruning the Deepseek-MoE-16B with different strategies.
 4.4 Pruning the Qwen Model
@@
 Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
-Table 3: Evaluation results of different pruning setting with ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
+Table 3: Evaluation results when ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
 # of samples MMLU BoolQ OpenBookQA RTE Avg.
-128 0.615 0.856 0.328 0.686 0.621
-256 0.614 0.854 0.332 0.697 0.624
-512 0.617 0.854 0.334 0.704 0.627
-2,048 0.612 0.856 0.322 0.675 0.616
+128 61.5 85.6 32.8 68.6 62.1
+256 61.4 85.4 33.2 69.7 62.4
+512 61.7 85.4 33.4 70.4 62.7
+2,048 61.2 85.6 32.2 67.5 61.6
 4.6 Ablation Study and Discussion
@@
 The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result.
-We provide the full enumeration results in section A.1.
-On the number of samples.
-We conduct experiments to study the effect of the used sample in data-centric pruning methods.
-As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$.
-It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
-We attribute this phenomenon to the use of mean operation in our expert similarity estimation.
-On the one hand, the mean operation keeps the feature dimension for efficient pairwise distance computation.
-On the other hand, it naturally downgrades the precision of function similarity computation by “round-off” error compared with other per sample-based distance metrics like CKA used in section 3.2.
-On merging the routing policy.
-One relevant work to ours is [21], which identifies and merges similar experts based on routing policy, setting the routing policy of the merged expert to that of the most frequently visited expert.
-While we identify similar experts based on expert knowledge, such as weights and features, we merge not only the experts but also the routing policies.
-To demonstrate the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies.
-As shown in table 4, this results in performance degradation across all tasks and different numbers of experts, highlighting the crucial role of simultaneously merging the routers and experts.
-Table 4: Evaluation results of our augmentation-based pruning method without merging the routers.
+We provide the full enumeration results in appendix C.
+On the used samples.
+We conduct experiments to study the effect of both sample size and augmentation in data-centric pruning methods.
+As shown in table 3, increasing the number of samples from $128$ to $512$ without augmentation leads to a noticeable performance improvement of up to $0.6\%$.
+However, when the sample size is further increased to $2,048$, the average performance drops significantly to $61.6\%$.
+In contrast, when comparing this result with table 1, where $128$ samples were used with augmentation, we observe an average performance improvement of $0.3\%$.
+This demonstrates the importance of both sample size and augmentation in enhancing generalization during the pruning process, as discussed in section 3.4.
+We attribute the performance drop with large sample sizes to the mean operation in our expert similarity estimation.
+While the mean preserves the feature dimension for efficient pairwise distance calculations, it reduces the precision of similarity estimates due to "round-off" errors, compared to per-sample metrics like CKA used in section 3.2.
+On merging the routing policy.The motivation of our work is to diversify expert knowledge by merging similar experts.
+A key distinction of our approach is the merging of routing policies, which potentially directs more tokens to the resulting merged expert.
+To highlight the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies.
+As shown in table 4, compared with the results in table 1 and table 2,this results in performance degradation across all tasks and different numbers of experts, underscoring the crucial role of simultaneously merging both routers and experts.
+Table 4: Evaluation results of our data-centric pruning method without merging the routers.
 # of experts MMLU BoolQ OpenBookQA RTE Avg.
-4 0.491 0.833 0.274 0.664 0.565
-6 0.604 0.847 0.312 0.675 0.609
+4 49.1 83.3 27.4 66.4 56.5
+6 60.4 84.7 31.2 67.5 60.9
 5 Conclusion and Future Work
-In this paper, we address the practical problem of reducing the number of experts in MoE layers to accelerate the inference and fine-tuning process.
+In this paper, we work on the task-agnostic pruning of sparse MoEs.
 We propose discovering similar experts at the feature level and then merging them in the weight space for MoE pruning while preserving as much original expert knowledge as possible.
 This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
-However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers.
+However, we didn’t account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers.
 Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
-6 Acknowledgments
-We would like to thank Lifeng Li from Microsoft for setting up GPU machines.
-Appendix A Evaluation on the expert similarity
-Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B.
-The results are respectively depicted in fig.
-A1 and fig.
-A2.
-We summarize the observation as follows,
-  • Most MoE layers in the two Mixtral models contain significant expert redundancy.
-  • The most redundant MoE layers are located in the first and last several layers, while the experts in the intermediate MoE layers learn more diverse features.
-Figure A1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
-Figure A2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.
-A.1 Enumeration on the expert pruning
-We present the full greedy search result on Mixtral-8x7B in fig.
-A3.
-In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task.
-For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
-Although dropping most of the combinations on different layers only leads to a minor performance drop, we can notice that it could cause the model to crash when the fourth expert in layer 1 is involved during the pruning process.
-Figure A3: Enumeration on dropping two experts.
-A.2 Results on different strategies for pruning
-While we report the best performance using our data-centric and model-centric strategies to prune the MoE models, we detail more results on pruning the Mixtral-8x7B and Mixtral-8x22B.
-For model centric, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts.
-We uniformly merge the experts and their routers for model pruning.
-For data centric, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), or only maintain the expert with the maximum visiting frequency (Max).
-From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods.
-This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
-2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance.
-Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
-Table A1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer.
+6 Limitations
+Several unexplored questions remain in our project.
+First, we designed various strategies to prune the MoE, and we observed that different models require different strategies to achieve optimal post-pruning performance.
+It remains unclear what causes these performance differences across strategies.
+Second, while the learning strategy at the merging step can bring slightly performance improvement, the cost is also large.
+The question of how to efficiently find the optimal merging coefficients remains.
+Third, in our work, we prune the same experts across different MoE layers, despite each layer having varying levels of redundancy.
+A key question remains: how can we push MoE compression to its limits while maintaining acceptable performance?
+Appendix A Experiment details
+We conducted our experiments using four 80GB NVIDIA A100 GPUs.
+In our learning strategy, we used SGD as the optimizer to learn the coefficients for expert merging, with the learning rate set to $1\times 10^{-3}$.
+The coefficients were initialized as identity matrices and optimized over $50$ epochs.
+For baseline methods and models, all usage and distribution comply with the terms of their license, i.e., Mixtral (Apache License 2.0), Deepseek (MIT license), Qwen (Tongyi Qianwen license), lm evaluation harness (MIT license).
+Appendix B Evaluation on the expert similarity
+Table B1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer.
 We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq.
@@
 Method humanities social science stem Other
-Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
-Model-centric Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
-Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
+Mixtral-8x7B 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
+Model-centric Vectorized 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
+Surrogate 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
 […]
 Max 56.4/47.6 71.9/58.6 52.0/42.9 66.9/55.7 85.1/82.8 35.2/28.8 70.4/66.8 62.6/53.7
-Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
-Model-centric Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
-Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
+Uniform 56.2/47.8 72.7/57.2 52.1/42.7 68.1/54.0 85.6/83.2 32.8/28.6 68.6/66.5 62.3/54.3
+Mixtral-8x22B 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
+Model-centric Vectorized 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
+Surrogate 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
 […]
-Max $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
+Max 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
+Uniform 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
+Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B.
+The results are respectively depicted in fig.
+D1 and fig.
+D2.
+We summarize the observation as follows,
+  • Most MoE layers in the two Mixtral models contain significant expert redundancy.
+  • The most redundant MoE layers are located in the first and last several layers, while the experts in the intermediate MoE layers learn more diverse features.
+Appendix C Enumeration on the expert pruning
+We present the full greedy search result on Mixtral-8x7B in fig.
+D3.
+In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task.
+For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
+Although dropping most of the combinations on different layers only leads to a minor performance drop, we can notice that it could cause the model to crash when the fourth expert in layer 1 is involved during the pruning process.
+Appendix D Results on different strategies for pruning
+While we report the best performance using our data-centric and model-centric strategies to prune the MoE models, we detail more results on pruning the Mixtral-8x7B and Mixtral-8x22B.
+  1. 1.
+    For model-centric stratgies, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts.
+We uniformly merge the experts and their routers for model pruning.
+In other words, the model-centric strategies differ at the discovery stage.
+  2. 2.
+    For data-centric strategy, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), only maintain the expert with the maximum visiting frequency (Max), or uniformly merging different experts (Uniform).
+Thus, in this experiment, the data-centric strategies differ at the merging stage.
+From the results of our strategies, we can observe the following:
+  1. 1.
+    Both the Vectorized and Surrogate strategies surpass all other data-involved pruning baseline methods.
+This suggests that weights already encode valuable data information, which can be utilized to group experts for pruning.
+  2. 2.
+    When pruning relatively small models (Mixtral-8x7B), the inclusion of data in the pruning process improves candidate selection for merging and pruning, leading to better performance compared to using only model weights.
+While the learning strategy offers a slight improvement in average performance, it comes at a higher computational cost compared to using uniform coefficients for merging.
+  3. 3.
+    For larger models (Mixtral-8x22B), the model-centric method outperforms the data-centric method.
+We argue that this is due to overfitting to the small calibration dataset during pruning, especially when dealing with a large number of parameters.
+The small calibration dataset cannot approximate the distribution of the pre-training dataset effectively.
+This is evident from the significant performance drop observed when using the Max strategy for pruning Mixtral-8x22B.
+Figure D1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
+Figure D2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.
+Figure D3: Enumeration on dropping two experts.
+Appendix E Full results on pruning Deepseek and Qwen
+We present the full results of pruning Deepseek-MoE-16 and Qwen2-57B-14A in fig.
+E4 and fig.
+E5.
+It is evident that our proposed data-centric and model-centric strategies outperform all baseline methods in most test cases.
+Additionally, when examining different evaluation tasks, the results on MMLU show a more reliable and consistent trend as the number of pruned experts increases.
+However, for smaller tasks such as the RTE dataset, we observe some randomness in the evaluation results due to the limited dataset size.We did not include the enumeration-based method in our comparison, as it is time-consuming and difficult to complete within a limited timeframe, especially when the number of experts is large in these models.
+Figure E4: Results on pruning Deepseek-MoE-16B.
+Figure E5: Results on pruning Qwen2-57B-14A.

### v4

@@
 Abstract
-By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost.
-However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real-world settings.
+In this work, we address the memory overhead of deploying Mixture-of-Experts (MoE) architectures in Large Language Models (LLMs).
+While MoE layers improve LLM performance without increasing inference costs, the ever-growing number of experts inflates memory requirements, hindering practical deployment.
 Our empirical study reveals that some experts encode redundant knowledge during pre-training.
@@
 The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
-We will release our code to facilitate future research.
 […]
@@
 in Sparse Mixture-of-Experts
-Zeliang Zhang1111This research was conducted during the author’s internship at Microsoft Research.
-Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
+thanks: Work done during the internship at Microsoft Research.Zeliang Zhang1  Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
 1 Introduction
@@
 (2024a) merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning.
-Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks.
+While effective, these methods depend on knowing the target tasks.
+In contrast, task-agnostic pruning methods that do not rely on task information are more appealing and useful in real-world applications because they can apply to both seen and unseen tasks.
+However, this is more challenging since there are no explicit task and data cues to guide which experts are redundant.
 He et al.
@@
 The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
-Our contribution can be summarized as follows,
+Our contributions are summarized as follows,
   1. 1.
@@
 In this paper, we work on pruning the sparse MoEs to reduce redundant experts in the task-agnostic setting, which enhances the computational and memory efficiency throughout the fine-tuning process, and scalability of deploying these models.
-3 Methodology
-Notations.
-We denote $F(\cdot;\Theta,W,K)$ as an MoE layer, where $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ indicates the parameter set for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, $W\in\mathbb{R}^{N\times d}$ indicates the routing weights for each expert and $K$ indicates the number of experts selected for each token.
-For a new token $x\in\mathbb{R}^{d}$, we first compute the routing logits to score the match between $x$ and the $n$-th expert:
-$p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}},$ (1)
-where $W_{t}\in\mathbb{R}^{d}$ is the routing policy of $t$-th expert.
-Then, we take the experts $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$ with top-$K$ logits $\{l_{n}\}_{n=i_{1}}^{i_{K}}$ as the selected target experts.
-Last, we compute the weighted average of the outputs from selected experts as the final output of the MoE layer,
-$y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$ (2)
+3 Method
+In this section, we present our approach for pruning experts within MoE layers in large language models.
+Our goal is to identify experts that share highly similar knowledge and merge them, thereby reducing the model size and improving efficiency without significantly degrading performance.
+3.1 Notation and Preliminaries
+We denote $F(\cdot;\Theta,W,K)$ as an MoE layer.
+Here, $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ is the set of parameters for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, where each expert is typically a feed-forward network (FFN).
+The routing matrix $W\in\mathbb{R}^{N\times d}$ determines which experts are selected for each token.
+K denotes how many top experts are chosen for each input token.
+For a token $x\in\mathbb{R}^{d}$, we first compute the routing logits to measure how well it matches each expert: $p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}}$.
+Then, we select top-$K$ experts, $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$, based on these logits.
+Finally, we compute the MoE layer’s output as the weighted combination of the chosen experts: $y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$
 3.2 Task Definition
-Previous works (Lu et al., 2024; He et al., 2024; Li et al., 2024a) have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks.
-The remaining question is which expert(s) should be pruned in the MoE layer.
-Specifically, for the $l$-th MoE layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ in the given LLM $\mathcal{M}(\cdot;\mathcal{F})$, we want to prune $r$ experts from the total of $N$, $l=1,2,...,L$.
-This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete.
-We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ towards the following optimization objective:
-$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y),$ (3)
-where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the new LLM with pruned MoE layer $\hat{\mathcal{F}}$, and $\mathcal{D}$ is any dataset under the task-agnostic setting.
-This task is particularly challenging, as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions.
-In this work, we propose to estimate the expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
+Previous works (Lu et al., 2024; He et al., 2024; Li et al., 2024a) have shown that pruning some experts in MoE layers can improve model efficiency in both inference and fine-tuning without causing model collapse.
+However, an open question remains: which experts should be removed?
+Formally, consider a LLM $\mathcal{M}(\cdot;\mathcal{F})$ comprsing $L$ MoE layers.
+Each MoE Layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ has $N$ experts.
+Our objetive is to only remain $r$ experts from each MoE layer:
+$\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{N-r}}\}$ (1)
+$\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{N-r}}\}$
+We aim to find these indices $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{N-r}\}$ that minimizes loss on a generic dataset $D$:
+$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y).$ (2)
+The search space $(C^{r}_{N})^{L}$ is enormous, making direct combinatorial search intractable.
+3.3 Measuring Expert Similarity with CKA
 Figure 2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under two kernel-based CKA criteria (Linear and RBF).
 A darker color indicates a greater similarity between experts.
-3.2 Expert Similarity in MoE layer
-Motivation: Intuitively, the more independent the experts are within the same MoE layer, the more diverse the representations learned by each expert.
-This diversity makes it harder to replace or drop any of the experts.
-Conversely, the redundancy of shared knowledge learned by different experts provides opportunities for model pruning.
-We employ the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) as the criteria to evaluate the similarity between experts in each MoE layer.
-Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
-$\rho_{ij}=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (4)
-where $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
-Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 (Raffel et al., 2020).
-We respectively use the linear kernel and RBF kernel in similarity computation of eq.
-4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
-As illustrated in Fig.
-2, the darker the cell, the more similar the two related experts.
-The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers.
-There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts.
-For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
+Our key insight is that experts exhibiting similar behaviors likely contain redundant knowledge. By identifying and pruning these redundant experts, we can simplify model and improve its efficiency.
+To quantify similarity, we use Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Davari et al., ; Smerkous et al., ) as the criteria to evaluate the similarity between experts in each MoE layer.
+Intuitively, CKA measures how similarly two experts transform a shared batch of inputs.
+For any two experts $f_{i}$ and $f_{j}$, given a batch of inputs $x=\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ is computed as follows,
+$\rho_{ij}=\text{CKA}(K^{i},K^{j})=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (3)
+where $K^{i}$ and $K^{j}$ are kernel matrices constructed from experts’ outputs on the same input batch $x$, and HSIC is the Hilbert-Schmidt Independence Criterion (Greenfeld and Shalit, 2020).
+The definition is as follows: $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
+Notably, all experts are provided with the same input batch—no token distribution by the router function is involved—allowing for a clearer representation of the distinct knowledge each expert acquires within the same layer.
+We evaluate the similarity between experts in Mixtral-8Bx7B using 32 randomly selected samples from the C4 pre-training dataset (Raffel et al., 2020).
+Following eq.
+3, we compute expert similarity with both a linear kernel (equivalent to using a dot product) and an RBF kernel.
+As shown in fig.
+2, darker cells indicate greater similarity between pairs of experts.
+Both kernel measures reveal similar patterns across layers, indicating that some experts are moderately to highly similar.
+This suggests the feasibility of pruning certain MoE layers.
+For example, in the first layer, experts 2 through 5 consistently show similarity scores above 0.7, suggesting that removing one of these experts would likely have minimal impact on overall model performance.
 Full evaluation results can be found in appendix B.
-3.3 Discover and Merge Similar Experts on Feature Space
+3.4 Discovering and Merging Similar Experts
+As shown in fig.
+3, our pruning strategy consists of two key steps to reduce the number of experts from $N$ to $r$: (1) identifying groups of similar experts and (2) merging each group into a single expert.
 Figure 3: We first leverage model or data-centric strategies to obtain the expert representation, then compute the pairwise distance to get the disparity matrix.
 Based on the expert similarity matrix, we can group similar experts with shared knowledge in the same cluster, which can be merged on the weight space for pruning.
-Having identified the existence of moderate expert similarity in each MoE layer, we propose pruning the experts within each MoE layer by merging experts with shared knowledge.
-We present an overview of our method in fig.
-3.
-Suppose we want to reduce the number of experts from $N$ to $r$, the details of our approach are outlined as follows.
 Discovering Similar Experts from the Expert Graph.
-We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ to group similar experts into the same clusters, where we denote $\mathcal{V}$ as the set of nodes and $\mathcal{E}$ as the connections between nodes.
-First, for each expert $f_{i}$, we compute its feature representation $\mathcal{M}(f_{i})$, where $\mathcal{M}(\cdot)$ is a specialized function to represent $f_{i}$.
-Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
-Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the pair-wise expert distance matrix $D$.
+We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A}\}$ to cluster similar experts, where $\mathcal{V}$ is the set of nodes (each representing an expert), $\mathcal{E}$ denotes the set of edges connecting experts that exhibit positive similarity, and $\mathcal{A}$ is a weight matrix that encodes these pairwise similarities.
+The procedures for constructing $\mathcal{A}$ is shown as follows.
+First, we represent each expert $f_{i}$ by its output embedding on a shared calibration dataset (data-centric) or by its own weights (model-centric) if data is unavailable, as $\mathcal{R}(f_{i})$.
+Second, we calculate the weight on each edge of $\mathcal{E}$ to represent the similarity between experts, i.e., $\mathcal{A}\left(\mathcal{E}\right)$, using CKA (data-centric) as eq.
+3 or mean squared error (model-centric)***More discussion is provided in Appendix G.
+Each element $\mathcal{A}\left(\mathcal{E}_{ij}\right)=\text{CKA}(\mathcal{R}(f_{i}),\mathcal{R}(f_{j}))$ reflects how similar experts $f_{i}$ and $f_{j}$ are in the transformed space $\mathcal{R}(f)$ as the previous step.
 Last, we split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$ to group similar experts.
@@
 This can be formulated as the follows,
-$\min\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}D_{jk}-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}D_{jk}\right),$ (5)
+$\max\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}\mathcal{A}\left(\mathcal{E}_{jk}\right)-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}\mathcal{A}\left(\mathcal{E}_{jk}\right)\right),$ (4)
 […]
-where The objective function seeks to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $D_{jk}$ based on expert $i$ and $j$.
-Merging Similar Experts with Routers.
-To diversity the experts and preserve the different knowledge learned by different models clustered in the same group (Wortsman et al., 2022; Wang et al., 2022), we merge the clusted experts with their routers on the weight space as follows,
+Our objective is to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $\mathcal{A}\left(\mathcal{E}_{ij}\right)$ based on expert $i$ and $j$.
+Merging Similar Experts To diversify the experts and preserve the different knowledge learned by different models clustered in the same group (Wortsman et al., 2022; Wang et al., 2022; Zhao et al., 2024), we merge the clustered experts with their routers on the weight space as follows,
 $\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},~{}~{}\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$ (5)
@@
 We update the MoE layer $F$ by respectively replacing all the experts grouped in $\mathcal{G}_{n}$ by a single FFN expert layer $f(\cdot;\hat{\theta}_{n})$ and corresponding routing weights $\hat{W}_{n}$, $n=1,2,...,N-r$.
-3.4 Practical Implementation
-On the discovery step. We propose two strategies for computing the disparity matrix $D$, including the data-centric strategy and model-centric strategy.
+The pseudo-code is provided in appendix G.
+3.5 Practical Considerations
+Computing the disparity matrix $\mathcal{A}(\mathcal{E})$.
+We propose two strategies for computing the disparity matrix $\mathcal{A}(\mathcal{E})$: a data-centric strategy and a model-centric strategy.
 For the data-centric strategy, since the full pre-training datasets for different models are large and inaccessible, we use the C4 dataset, a commonly used smaller pre-training subset that serves as an effective surrogate for capturing task-agnostic knowledge of experts Lu et al.
 (2024).
-Specifically, by disabling the router function, for the same input, we use the output of each expert as its expert representation $\mathcal{M}(\cdot)$.
+Specifically, by disabling the router function, for the same input, we use the output of each expert as its expert representation $\mathcal{R}(\cdot)$.
+Then, we compute the expert similarity and discover the similar experts by eq.
+4.
 To enhance generalization and mitigate overfitting to the selected samples during model pruning, we apply data augmentation by randomly mixing token embeddings during the representation computation at the discovery stage.
@@
 One is to leverage the vectorized weights directly ($\clubsuit$).
-The other is to leverage the local linearity of neural networks (Zhang and Wu, 2020) to compute the surrogate weight matrix ($\spadesuit$).
+The other is to leverage the local linearityIn short, the ’local linearity’ property of neural networks refers to the observation that, despite having multiple hidden layers and non-linear activation functions, neural networks exhibit linear behavior within a local region.
+This property motivates our design of a single-layer approximation using the surrogate weight.
+of neural networks (Zhang and Wu, 2020) to compute the surrogate weight matrix ($\spadesuit$).
 Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$.
@@
 Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance (see appendix D).
-On the merging step.
+The selection of distance function.
+While the CKA metric provides an accurate estimation of expert similarity, it can be both memory- and computation-intensive.
+As an alternative, cosine similarity may be used as an efficient approximation.
+This corresponds to computing CKA for a single sample (data-centric), replacing the $K_{i}$ and $K_{j}$ in eq.
+3 with the weights of the $i$-th and $j$-th experts, respectively (model-centric).
+In practical deployments, we also observe that mean squared error performs well across several models, including DeepSeek-MoE and Qwen-MoE.
+Discussion on two strategies.
+Ensuring distributional similarity between surrogate and real datasets is essential for the data-centric method to accurately estimate expert similarity; without the guarantee, its performance could be degraded significantly.
+While pre-trained data for many public MoEs, such as Mixtral, DeepSeek, and Qwen, is unavailable, our study, along with prior research, demonstrates that leveraging a small subset of the open-source dataset C4 can still achieve competitive performance.
+Alternatively, the model-centric approach presents a robust solution, eliminating the need for data access during pruning and, in some cases, achieving performance comparable to or exceeding that of data-centric methods.
+More discussion is provided in appendix H.
+Merging Strategies and Trade-offs.
 We consider three approaches in practice deciding $\alpha$ to merge similar grouped experts.
@@
 The Qwen model also has $28$ MoE layers with $64$ experts in each layer but will activate $8$ experts during the inference.
+We include more experimental details in appendix A.
 Pruning Methods.
@@
 The results are illustrated in fig.
-4***We only show the pruning results of the Deepseek and Qwen models on the most challenging MMLU task and the average performance across MMLU, BoolQ, OpenBookQA, and RTE.
+4We only show the pruning results of the Deepseek and Qwen models on the most challenging MMLU task and the average performance across MMLU, BoolQ, OpenBookQA, and RTE.
 Full results can be found in supplementary..
@@
 We prune the experts from $64$ experts to $48$ experts in each layer and evaluate the performance on different tasks.
-The results are reported in fig.
-5.
-When pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method.
-On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the count-guided pruning strategy.
-4.5 Empirical Analysis on Expert Hints
-While expert pruning brings a significant speedup by reducing the parameters, what changes are introduced to different MoE layers?
-To answer this question, we analyze the change of expert hints on the calibration dataset C4.
-Figure 6: Statistics of the visiting frequency for all experts in different MoE layers.
-The statistic results are shown in fig.
-6.
-We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints.
-In contrast, pruning differentiates the hints of experts.
-Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints.
-Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
+Reported in fig.
+5, when pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method.
+On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the runner-up one.
+Performance discrepancy between Deepseek and Qwen.
+The difference in performance appears to stem from variations in model architectures and training strategies.
+Specifically, Deepseek-MoE-16B uses shared experts to learn a broad range of knowledge, which can lead to highly diverse expert functions.
+Consequently, identifying similar experts for pruning becomes challenging, and removing them risks losing unique knowledge—resulting in notable performance drops.
+In contrast, Qwen2-57B-14A initializes its expert weights from a pre-trained dense model, increasing expert similarity post-training.
+Meanwhile, Deepseek-MoE-16B’s experts are randomly initialized and trained from scratch, resulting in fewer interchangeable experts.
+Pruning such independently specialized experts leads to a more pronounced performance decline.
 Table 3: Evaluation results when ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
@@
 512 61.7 85.4 33.4 70.4 62.7
-2,048 61.2 85.6 32.2 67.5 61.6
 4.5 Ablation Study and Discussion
@@
 To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method.
-Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer.
+We enumerate all possible combinations of removing two experts in each layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer.
 We then merge these results to prune the model across all layers.
 The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result.
-We provide the full enumeration results in appendix C.
+Full results are shown in appendix C.
 On the used samples.
-We conduct experiments to study the effect of both sample size and augmentation in data-centric pruning methods.
-As shown in table 3, increasing the number of samples from $128$ to $512$ without augmentation leads to a noticeable performance improvement of up to $0.6\%$.
-However, when the sample size is further increased to $2,048$, the average performance drops significantly to $61.6\%$.
-In contrast, when comparing this result with table 1, where $128$ samples were used with augmentation, we observe an average performance improvement of $0.3\%$.
+We conduct experiments to study the effect of both sample size used for the calibration dataset and augmentation in data-centric pruning methods.
+As shown in table 3, increasing the number of samples from $128$ to $512$ without augmentation leads to a noticeable performance improvement of up to $0.6\%$.In contrast, when comparing this result with table 1, where $128$ samples were used with augmentation, we observe an average performance improvement of $0.3\%$.
 This demonstrates the importance of both sample size and augmentation in enhancing generalization during the pruning process, as discussed in section 3.5.
-We attribute the performance drop with large sample sizes to the mean operation in our expert similarity estimation.
-While the mean preserves the feature dimension for efficient pairwise distance calculations, it reduces the precision of similarity estimates due to "round-off" errors, compared to per-sample metrics like CKA used in section 3.2.
 On merging the routing policy.The motivation of our work is to diversify expert knowledge by merging similar experts.
@@
 6 60.4 84.7 31.2 67.5 60.9
-5 Conclusion and Future Work
+Efficiency improvement after MoE pruning To demonstrate the benefits of MoE pruning, we provide a statistical efficiency analysis for Mixtral-8x7B and DeepSeek-MoE-16B.
+We set the batch size to 8 and evaluate the performance on the MMLU task using the lm_eval library (Gao et al., 2021).
+Other settings are consistent with the previous experiment.
+As shown in table 5, pruning two experts from each MoE layer in the Mixtral model achieves a 1.17× speedup and a 16.8% reduction in GPU memory usage.
+Pruning 16 experts from each MoE layer in the Deepseek model reduces inference time by 15.5% and GPU memory usage by 14.1%.
+These results highlight the computational benefits of expert pruning across different model scales.
+More analysis can be found in appendix F.
+Table 5: Computation cost of evaluating Mixtral-8x7B and DeepSeek-MoE-16B on the MMLU task. Experiment settings are consistent with our paper.
+Model # Experts Time (s) Mem (GB)
+Mixtral-8x7B 8 (Ori.) 281 125
+6 241 104
+4 223 83
+[…]
+60 446 61
+56 429 59
+52 413 57
+48 386 55
+5 Conclusion
 In this paper, we work on the task-agnostic pruning of sparse MoEs.
@@
 This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
-However, we didn’t account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers.
-Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
 6 Limitations
@@
 We conducted our experiments using four 80GB NVIDIA A100 GPUs.
-In our learning strategy, we used SGD as the optimizer to learn the coefficients for expert merging, with the learning rate set to $1\times 10^{-3}$.
-The coefficients were initialized as identity matrices and optimized over $50$ epochs.
+In our learning strategy for expert merging, we used SGD as the optimizer to learn the coefficients for expert merging, with the learning rate set to $1\times 10^{-3}$.
+We randomly sample $128$ samples from the calibration set (C4) and partition them into a $3:1$ ratio as the training and evaluation sets, respectively.
+The coefficients were initialized as identity matrices and optimized over $50$ epochs until they converged.
 For baseline methods and models, all usage and distribution comply with the terms of their license, i.e., Mixtral (Apache License 2.0), Deepseek (MIT license), Qwen (Tongyi Qianwen license), lm evaluation harness (MIT license).
+We use Copilot to help with debugging and coding.
 Appendix B Evaluation on the expert similarity
@@
 We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq.
-7; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency.
+6; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency.
 The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
@@
 Figure E5: Results on pruning Qwen2-57B-14A.
+Appendix F Empirical Analysis on Expert Hints
+We analyze the change of expert hints on the calibration dataset C4, where the expert hint refers to the visiting frequency of the expert on the calibration set.
+Figure F6: Statistics of the visiting frequency for all experts in different MoE layers.
+The statistic results are shown in fig.
+F6.
+We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints.
+In contrast, pruning differentiates the hints of experts.
+Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints.
+Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
+Appendix G Pseudo-code implementation
+Algorithm 1 Pruning algorithm for given MoE layer
+1:The set of experts =$\mathcal{F}=\{f_{1},f_{2},...,f_{N}\}$ in the given MoE layer to prune, each expert $f_{i}$ consisting of three layers parameterized by $\theta^{i}_{1},\theta^{i}_{2},\theta^{i}_{3}$, respectively, the number of experts to reserve $r$, the calibration dataset $\mathcal{D}$, strategy s.
+2:Pruned MoE layer with reduced experts $\{f_{1},f_{2},...,f_{r}\}$.
+3:function MoE pruning(MoE layer $\mathcal{F}$, number of experts $r$ to reserve)
+4:     Initialize $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A}\}$.
+5:     for each $i$ in $N$ do
+6:         if s==Data-centric then
+7:              $\mathcal{R}(f_{i})=f_{i}(\text{Mixup}(\mathcal{D}))$.
+8:         else if s==Vectorized weights then
+9:              $\mathcal{R}(f_{i})=\{\theta^{i}_{1},\theta^{i}_{2},\theta^{i}_{3}\}$
+10:         else if s==Surrogate weights then
+11:              $\mathcal{R}(f_{i})=\theta^{i}_{2}(\theta^{i}_{1}\cdot\theta^{i}_{3}$)
+12:     for each $i$ in $N$ do
+13:         for each $j$ in $N$ do
+14:              $\mathcal{A}(\mathcal{E}_{ij})=\text{CKA}(\mathcal{R}(f_{i}),\mathcal{R}(f_{j}))$
+15:     Optimize eq. 4 on $\mathcal{G}$ to find $r$ subgroup of experts.
+16:     Merge experts with their routers clustered within the same subgroup based on eq. 5.
+17:     return Pruned $\mathcal{F}$
+Solving eq.
+4 is an NP-hard problem.
+To obtain an approximate solution, two methods can be employed.
+One approach is to use spectral clustering (Ng et al., 2001), while the other is a greedy strategy that iteratively merges the pair of experts with the highest similarity to construct the pruned expert graph.
+Although both methods yield comparable performance, the greedy approach is significantly more efficient, particularly when dealing with large language models that have many experts per MoE layer.
+In practice, for data-centric method, we also observe that using cosine similarity performs on par with centered kernel alignment (CKA), while being more computationally efficient for measuring expert similarity.
+At the same time, the mean-squared error also works well, which we adopt in the model-centric method.
+Appendix H Discussion on data- and model-centric methods
+The data-centric method is a natural extension of the paper’s starting point.
+However, its implementation requires access to a small portion of the pre-training dataset or a surrogate dataset with a similar distribution, which may pose practical challenges.
+On the one hand, achieving good distributional similarity between the surrogate and real datasets is indeed crucial for the success of the data-centric method.
+Without such similarity, the method can fail entirely.
+In our trials, we tried using random inputs to identify similar experts and prune the MoE model, which led to the crashed performance of the pruned experts.
+On the other hand, while surrogate datasets are generally simpler than real-world datasets, they still capture critical characteristics learned by different experts.
+Notably, no pre-trained data is publicly available for the models we study, including Mixtral, DeepSeek, and Qwen.
+However, leveraging a small portion of the open-source pre-trained dataset C4 still yields comparable performance on diverse benchmarks, as shown in our results.
+In contrast, the model-centric method identifies similar experts directly by analyzing the model’s weights, bypassing the need for any data.
+This approach builds on the observation that training on the same dataset produces similar weight distributions, which can serve as a surrogate for predicting outputs given the same inputs.
+In the context of MoEs, when the gating mechanism consistently routes similar tokens to a group of experts, those experts tend to exhibit similar weights.
+Therefore, the model-centric strategy aligns with the paper’s initial data-dependent premise by offering a complementary perspective on expert similarity without relying on external datasets.